Solve the equation y'=x(1+y^2) by using variable separable
1
Expert's answer
2019-10-11T10:20:26-0400

The differential equation is



(1) "\frac {dy}{dx} =x(1+y^2)."



We saw that this is a separable equation, and can be written as


"\frac {dy}{1+y^2} =x{dx}."



Let’s take the integrals from both parts of the equation:



"\int \frac {dy}{1+y^2} = \int x dx,"


so "\,"


"\arctan(y) = \frac {1}{2}x^2 +C, C=Const.\\"


Therefore, the general solution of a differential equation (1):


"y = \tan (\frac {1}{2}x^2 +C)."